Question with last attempt is displayed for your review only Test the claim that the proportion of people who own cats is significantly different than 10% at the 0.05 significance level. The null and alternative hypothesis would be: Ho: μ 0.1 Ho:p> 0.1 Ho:p-0.1 Ho: μ < 0.1 Ho: p < 0.1 Ho : μ> 0.1 Hing 0.1 Hi p < 0.1 Hi : p+ 0.1 Hing 0.1 14:P> 0.1 HO 0.1 The test is: two-tailed right-tailed left-tailed Based on a sample of 800 people, 6% owned cats The test statistic is: D The p-value is: 00 (to 2 decimals) (to 2 decimals) Based on this we: O Fail to reject the null hypothesis Reject the null hypothesis
Answer #1

(a)

Correct option:

H0: p = 0.1

H1: p 0.1

(b)

Correct option:

two - tailed

(c)

n = sample size = 800

P = population proportion = 0.1

Q= 1- p = 0.9

SE =

p = Sample proportion = 0.06

test statistic is:

Z =(0.06 - 0.1)/0.0106 = - 3.77

(d)

Table of Area Under Standard Normal Curve gives area = 0.4999

So,

P value = (0.5 - 0.4999) X 2 = 0.00

(e)

Correct option:

Reject the null hypothesis